OpenStudy (anonymous):

Find the equation of the tangent line to the following curve at the indicated point (20, 5). y^2=x^2/xy-84. y(x)=?

OpenStudy (anonymous):

find slope dy/dx and use y-y1=slope(x-x1)

OpenStudy (anonymous):

would i find the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y/2y^3+x would this be it?

OpenStudy (anonymous):

have you taken calc I

OpenStudy (anonymous):

taking it

OpenStudy (anonymous):

I dont understand this thou and the hw is due soon so im trying to finish it.

OpenStudy (anonymous):

Use implicit differentiation to find dy/dx of the given function. Substitute the x-value into the derrivative to find the gradient/slope of the tangent. Then use y-5=m(x-20) to find the equation of the tangent. m=gradient

OpenStudy (anonymous):

y/2y^3+x is this it?

OpenStudy (rogue):

\[y^2 = \frac {x^2}{xy-84}\]\[xy^3 - 84y = x^2\]Now use implicit differentiation\[\frac {d}{dx} \left[ xy^3 - 84y = x^2 \right]\]\[\frac {d}{dx} \left[ xy^3 \right] - 84 \frac {dy}{dx} = 2x\]\[y^3 + 3xy^2 \frac {dy}{dx} - 84 \frac {dy}{dx} = 2x\]\[\frac {dy}{dx} (3xy^2 - 84) = 2x - y^3\]\[\frac {dy}{dx} = \frac {2x - y^3}{3xy^2 - 84}\]

OpenStudy (anonymous):

To do the solution I need to understand your question clearly. is the function y^2=x^2/(xy-84) or its y^2=x^2/(xy) -84 ?

OpenStudy (anonymous):

Rogue's simplification of the given function is incorrect. Should have been xy^3 - 84y^2 = x^2. the error is on 84y, it should be 84y^2 i:e y should be squared

OpenStudy (rogue):

:3 damn, those little mistakes...

OpenStudy (anonymous):

-y^3+2x/3y(xy-56)

OpenStudy (anonymous):

that good?

OpenStudy (rogue):

\[\frac {d}{dx} \left[ xy^3 - 84y^2 = x^2 \right]\]\[y^3 + 3xy^2 \frac {dy}{dx} - 168y \frac {dy}{dx} = 2x\]\[\frac {dy}{dx}(3xy^2 - 168y) = 2x - y^3\]\[\frac {dy}{dx} = \frac {2x - y^3}{3y(xy - 54)}\]

OpenStudy (rogue):

Actually, mansukh's thing is more right, lol, I made a mistake dividing 168 by 3... epic fail.

OpenStudy (rogue):

\[\frac {dy}{dx} = \frac {2x - y^3}{3y(xy - 56)}\]

OpenStudy (anonymous):

lol yeah... now for the next step Mansukh, you need to plug in your given x and y values into the derrivative equation to get the gradient of the tangent

OpenStudy (anonymous):

then after that you can use straight line equation y=mx+c to get an equation of the tangent

OpenStudy (anonymous):

the one that rouge put up? 2(20)-5^3/3(5)((20)(5)-56)=-9126.6666

OpenStudy (anonymous):

Its actually -0.1287878788. Maybe you made an error on your calculator. Can you re-check

OpenStudy (anonymous):

this is what i got 7.57576-0.128788 x i put it into the site and it wasnt rite =(

OpenStudy (anonymous):

the questions asking for what dy/dx=

OpenStudy (anonymous):

Are you ok now, or you still do not understand?

OpenStudy (anonymous):

it says the answers wrong

OpenStudy (anonymous):

There's gotta be something wrong in the given equation because the procedure and the final answer is correct. I just double checked now

OpenStudy (anonymous):

oh well =(

OpenStudy (rogue):

\[Y_{tangent} = \frac {-17}{132}(x-20) + 5 = \frac {250}{33} - \frac {17x}{132} \approx 7.576 - 0.129x\]Da website sucks.